OpenStudy (anonymous):

True or false: The average of the product of two functions on an interval is the product of the average values of the two functions on the interval

OpenStudy (anonymous):

No, because \[\int\limits f(x)g(x)\neq \int\limits f(x)\times \int\limits g(x)\]

OpenStudy (anonymous):

false that is. i think.